OpenStudy (anonymous):

A parabolic radar antenna has the shape given by y^2=8x,(units are feet).For maximum concetration of reflected energy, the antenna element is always placed at the focus of the parabolic antenna reflector.Suppose the antena is placed on a tower such that the center of reflector is 25ft above the ground . When an aircraft is tracked,the antena is pointed skyward at angle of 45.How high abowe the ground will the antenna element be when antena is in that position,and what equation describes the shape of the antenna?

OpenStudy (anonymous):

The length of the focus should be on the x-axis and an integer. When you tilt that length 45 degrees. the y-value can be found by taking the sin(45) = y/f, where f = focus length. \[y = (f *\sqrt{2})/2\]. 25 + y is the final height, but the key is finding the focus, which I need to look at again, but is possible pretty easily. the sin of 45 is sqrt 2/2 so that's how you getthe y valuein the noted equation.

OpenStudy (anonymous):

the eq of the antenna could any eq that allows for that focus length to be the apex of a structure from the center of the parabolic eq.

OpenStudy (anonymous):

Possible answers; 1) 25.7ft , x^2-2xy+y^2-8*2^(1/2)*x-8*2^(1/2)y=0 2) 25.7ft, x^2+2xy+y^2-8*2^(1/2)*x+8*2^(1/2)y=0 3) 26.4ft x^2-2xy+y^2-8*2^(1/2)*x-8*2^(1/2)y=0

myininaya (myininaya):

alma?